An ac generator supplies an rms voltage of 5.00 V to an RL circuit. At a frequency of 20.0 kHz the rms current in the circuit is 45.0 mA; at a frequency of 26.0 kHz the rms current is 40.0 mA.

What is the values of L in this circuit?

What is the values of R in this circuit?
Answer #1

as we know

Z = V/i

so for case 1

R^2 + (2 pi * 20* 1000* L) ^2 = (5/0.045)^2

R^2 + 1.58* 10^10 L^2 = 12345.67 ...(i)

similarly, for case 2

R^2 + ( 2 * 3.14* 26000* L)^2 = (5 / 0.04)^2

R^2 + 2.67* 10^10 L^2 = 15625... (ii)

solving (i) and (ii)

L = 5.485* 10^-4 H

R = 87.2 ohm

=====

comment before rate in case any doubt, will reply for sure.. goodluck